# Generative Ghosts: Anticipating Benefits and Risks of AI Afterlives

### v2

@@ Introduction
We introduce the notion of a generative ghost, an agentic, AI-powered representation of a deceased individual. Generative ghosts differ from other technology-mediated representations of the deceased (such as memorialized social media accounts) in that they provide more than static or interactive access to content about the deceased produced during their lifetime (e.g., text, imagery, audio, video, or other media). Rather, they additionally are capable of generating novel content and interactions based upon the deceased’s data, potentially mediated by additional metadata such as personality questionnaires (e.g, (NORC at the University of Chicago n.d.)), prompts (Zamfirescu-Pereira et al. 2023), constitutions (Bai et al. 2022), or other rules set forth by the generative ghost’s creator. In addition to producing novel content, generative ghosts may also possess agentic (Park et al. 2023; Shavit et al. 2023) capabilities (i.e., the ability to autonomously execute complex sequences of actions with side effects in either the digital or physical realms).
@@
In the rest of this section, we expand on our definition of generative ghosts by discussing the design space of potential instantiations of this concept. While all generative ghosts share the capability to retrieve pre-existing multimodal content and generative novel multimodal content representing a deceased individual (including through the execution of agentic interactions), we argue that several key dimensions may impact the capabilities, perceptions, and societal impacts of such systems, including provenance, deployment timeline, anthropomorphism paradigm, multiplicity, cutoff date, embodiment, and representee type. As we detail in the Discussion, interrelationships between these dimensions are common. However, we present them separately here for the sake of clarity.
Provenance: First-Party vs. Third-Party Ghosts
@@
The provenance of a generative ghost refers to who created it (or authorized its creation). A first-party generative ghost is one created by the individual being represented (i.e., perhaps via end-of-life planning or legacy crafting practices). In contrast, a third-party generative ghost is not created by the representee. Third-party generative ghosts might be created by parties with a direct personal connection to the deceased (e.g., family, friends), by parties with a fiscal connection to the deceased (e.g., employers, estates), or by unconnected parties. Additionally, third-party ghosts may be authorized or unauthorized by the representee. Authorized third-party generative ghosts, for example, might be created with consent (e.g., via the deceased’s will). Meanwhile, unauthorized ghosts created by unconnected third parties are most likely in the case of public figures such as historical figures or contemporary celebrities; indeed, we already see early attempts at creating third-party generative ghosts of historical figures for entertainment and educational purposes by companies such as character.ai (Metz 2023) and nonprofits such as Khan Academy (Grant 2023).
Deployment Timeline: Pre-Mortem vs. Post-Mortem
@@
The deployment timeline of a generative ghost indicates whether the system’s initial deployment was pre-mortem or post-mortem. While some generative ghosts will be deployed post-mortem with the explicit purpose of memorializing the dead, many generative ghosts will likely start off as agents that serve some pre-mortem functionality (e.g., responding to personal emails or accomplishing other tasks on behalf of the representee, in character as if the representee had executed the task themself) – what we have termed generative clones – only to transition to ghosts after the death of their representee. Pre-mortem deployments might be desirable as they would allow an individual the opportunity to tune the behavior and/or capabilities of a first-party generative ghost to their liking, perhaps as part of an end-of-life planning process. Likewise, we imagine that generative clones might be designed with mortality in mind, potentially with attendant modifications to their behavior and/or capabilities once they become ghosts.
Anthropomorphism Paradigm: Reincarnations vs. Representations
@@
The anthropomorphism paradigm of a generative ghost refers to a subtle, but potentially impactful, interface metaphor choice: whether a generative ghost presents itself as a reincarnation of the deceased individual or as a representation of that individual. Generative ghosts employing a reincarnation metaphor would engage with end-users as if they were the deceased individual. In contrast, generative ghosts employing a representation metaphor would engage with end-users as if they are a representation of the deceased, but are not the deceased themselves. Design choices that might impact perceptions of whether a generative ghost is a reincarnation or a representation include whether a generative ghost uses first-person pronouns, whether it uses the present or past tense when discussing its representee, whether it uses the name of its representee or a different nomenclature (e.g., the use of “Fredbot” in (Kurzweil 2023)), whether it is allowed to make statements that assert it is alive, possesses a soul, etc.
Multiplicity: Single Ghost vs. Multiple Ghosts
@@
The multiplicity design dimension refers to whether an individual is represented by a single generative ghost or by multiple distinct ghosts. Multiple ghosts might arise intentionally (i.e., a single creator develops a set of generative ghosts with different behaviors, capabilities, and/or audiences in mind). Multiple ghosts might enable people to customize different ghosts for different audiences in order to avoid forms of context collapse (Davis and Jurgenson 2014; Litt 2012; Marwick and Boyd 2011). Multiple ghosts might also arise unintentionally due to multiple third parties creating generative ghosts for a single individual, or perhaps due to post-mortem identity theft or other novel cybercrimes.
Cutoff Date: Static vs. Evolving
@@
Embodiment: Physical vs. Virtual Ghosts
The embodiment dimension refers to whether a generative ghost has a physical embodiment. Such embodiments might be physical in a literal sense (e.g., robotics) or might be embodiments in rich digital media (e.g., avatars in mixed reality environments). In contrast, purely virtual generative ghosts would lack embodiment (e.g., perhaps available only as a chatbot). In addition to technical and/or cost constraints, there may be additional reasons to opt for lower-fidelity (i.e., virtual-only) embodiments, including potential ethical or psychological concerns related to physical embodiments (i.e., the strong anthropomorphic cues of embodied generative ghosts might elicit a reincarnation rather than representation metaphor, which might be undesirable for certain users or contexts).
Representee Type: Human vs. Non-Human Legacies
@@
Anticipating Benefits and Risks of Generative Ghosts
@@
Generative Ghosts: Potential Benefits
@@
Potential Benefits for Representees
Individuals may choose to create first-party generative ghosts as part of their end-of-life planning for several reasons. On a personal level, people may feel comforted by or even feel hopeful in the idea of a “digital afterlife.” Prior work on digital legacy has documented people’s motivations to capture their life stories (Thomas and Briggs 2014), major life events (Jamison-Powell et al. 2016), and pass on values to future generations (Gulotta, Kelliher, and Forlizzi 2017). Posthumous, personalized AI representations may likewise be seen as a way for people to ensure that they are remembered, whether by family, friends, or perhaps the world at large.
@@
Potential Benefits for the Bereaved
Interacting with a generative ghost may provide emotional support to close friends and family of the deceased. Prior research has considered the impact of online memorials on the bereaved (e.g. (Walter et al. 2012)), often responding to concerns that such spaces might prolong grief. Yet scholars cite the benefits of online memorials that allow loved ones to maintain healthy “continuing bonds” (Klass, Silverman, and Nickman 1996) with the deceased (e.g., (Getty et al. 2011; Brubaker and Hayes 2011; Klass, Silverman, and Nickman 1996); c.f., (DeGroot 2018)), often in a space where those grieving can gather (Carroll and Landry 2010; Brubaker, Hayes, and Dourish 2013).
As with griefbots more generally, discussions with a generative ghost might provide comfort or closure and support a continued feeling of closeness. Generative ghosts may support positive mental health, particularly if they enable the bereaved to maintain connections with the deceased while “making sense” of the loss (termed “accommodation” (Neimeyer and Thompson 2014)). Family members may find it comforting not only to interact with the generative ghost in the immediate post-mortem period, but to know that they might be able to interact with their deceased loved one at upcoming future life events such as weddings, the birth of a child, or other milestones during which they may wish to share updates with the generative ghost and/or hear advice, congratulations, or other event-appropriate commentary from their deceased loved one.
One-sided conversations with deceased loved ones, such as speaking graveside or writing to the deceased on a memorial, are common and well-recognized in the literature (DeGroot 2018; Brubaker and Hayes 2011; Brubaker, Hayes, and Dourish 2013) and are generally considered a part of healthy grieving (Klass, Silverman, and Nickman 1996). As summarized by DeGroot (DeGroot 2018), these conversations are cathartic and help the living confront their loss: “People essentially struggle to develop an identity where the deceased is now part of the past self and not the present self… and communicating with the deceased appears to ease the transition” (DeGroot 2018). However, it is worth noting that the dead do not typically respond. Rather, people imagine how the deceased would reply (DeGroot 2018). It unclear how responses from a generative ghost might impact bereavement.
@@
Potential Benefits for Society
@@
Generative Ghosts: Potential Risks
@@
While thoughtful implementations of generative ghosts have potential upsides for the representee, their loved ones, and society more broadly, there are also potential risks associated with this paradigm. By proactively anticipating these risks, we may be able to design technical and sociotechnical systems that avoid or mitigate them. Of course, as with any powerful new technology, it is impossible to fully anticipate apriori the many potential sociotechnial side-effects of its introduction; the emergence of “AI afterlives” may reshape society in complex ways beyond our current imagining (i.e., spurring the advent of new religious movements). Here, we discuss four broad categories of risk associated with generative ghosts: mental health risks, reputational risks, security risks, and socio-cultural risks.
Mental Health Risks
@@
Earlier we discussed ways in which generative ghosts might offer mental health benefits to mourners. However, ghosts may also pose mental health risks, including risks related to delayed accommodation, excessive attachment, information overload, anthropomorphism, deification, and second deaths. Studies of prior generations of AI technologies (and even of non-AI computing systems (Reeves and Nass 1996)) have found that people tend to anthropomorphize such systems and even develop attachments (Gabriel et al. 2024). We anticipate that without careful design generative ghosts might magnify these risks due to the increasing fidelity of generative AI technologies and the modeling of the agent on a specific individual.
@@
Interacting with a generative ghost may impact the bereaved’s ability to assimilate and accommodate (Payne, Joseph, and Tudway 2007; Neimeyer and Thompson 2014) their loss and integrate the reality of their loved one’s death (Klass, Silverman, and Nickman 1996). Delayed accommodation may negatively impact the bereaved’s mental health and quality of life by supporting loss-oriented experiences (e.g., active grieving, reminiscing while looking at old photos, etc.) at the expense of restorative-oriented experiences (e.g., the development of new relationships and life patterns) (Stroebe and Schut 1999). Stroebe and Schut argue that both loss- and restorative-oriented experiences are necessary, but that having agency around when and how one oscillates between the two is important – agency that could be impeded by the interactive design of generative ghosts (e.g., if a generative ghost were designed to use “push notifications” that might pull a mourner into an interaction rather than letting the mourner decide when and how they want to engage in remembrance and other grief-related activities).
@@
A related concern is the possibility of developing excessive attachment to generative ghosts, as indicated by spending a disproportionate amount of time interacting with representations of the deceased in lieu of interacting with the living. Already, there are indications that some individuals find AI-powered companions, such as those from Replika (replika.ai), to be highly compelling (Liang 2023); generative ghosts’ basis in real, beloved individuals rather than in fictional personas or historical figures may amplify this type of attachment.
@@
An additional mental health risk is that of second deaths or second losses, the experience of grieving a loved one again. [^1] Grief scholarship sometimes uses “secondary loss” to refer to experiences of grief that follow the “primary loss” of a loved one (e.g., loss of relationships, home, shared activities). Here, however, we focus on circumstances that might result in people engaging with their primary loss and grief a second time. In digital contexts, second deaths occur when data becomes unavailable either through technical obsolescence, deletion, or lack of access (Bassett 2022). In the case of generative ghosts, second death refers to the emotional harm that may result if a ghost ceases to exist. Second deaths may occur for myriad reasons, including economic (i.e., the company that maintains the generative ghost service goes out of business; survivors’ or the deceased’s estate can no longer afford maintenance fees), regulatory (i.e., a government outlaws generative ghosts, which may be particularly likely in theocracies where AI afterlives threaten established belief systems), technological (i.e., obsolete generative ghosts no longer run on future infrastructure), or security reasons (i.e., a hacker disables a generative ghost). While second loss is not unique to AI systems, the potential intensity of engagement that generative ghosts can support means that the loss experienced at the disabling of a generative ghost might be especially profound.
Reputational Risks
@@
Privacy risks are context-dependent (Nissenbaum 2009). For instance, it may be acceptable for a generative ghost to share sensitive information with some audiences, but not with others (e.g., perhaps it is acceptable for the ghost to discuss the deceased’s religious beliefs with their spouse, but not with their close work friend). The creation of multiple ghosts, each with different knowledge or abilities targeted toward different audiences, might mitigate privacy risks related to context collapse (Davis and Jurgenson 2014; Litt 2012; Marwick and Boyd 2011).
@@
Fidelity risks refer to challenges that might arise from both accurate and inaccurate information. Related to hallucination risks, ghosts providing inaccurate information can be a problem, particularly in historical, legal, and economic applications. Yet accurate information also comes with risks. As with human memories, personal and family stories change over the years, and most forms of media decay over time. However, digital media defaults towards persistence, which can impede the important roles that forgetting (Mayer-Schonberger 2009) and evolving memories can play (Pasupathi 2001). Generative ghosts might disrupt the natural tendency for loved ones to recall the deceased through “rose-colored glasses” (i.e., the tendency to focus on positive memories and attributes) by persistently reminding mourners of particular negative events or character traits.
Security Risks

### v3

@@ 3. Generative Ghosts: A Design Space
We introduce the notion of a generative ghost, an agentic, AI-powered representation of a deceased individual. Generative ghosts differ from other technology-mediated representations of the deceased (such as memorialized social media accounts) in that they provide more than static or interactive access to content about the deceased produced during their lifetime (e.g., text, imagery, audio, video, or other media). Rather, they additionally are capable of generating novel content and interactions based upon the deceased’s data, potentially mediated by additional metadata such as personality questionnaires (NORC at the University of Chicago, nd), prompts (Zamfirescu-Pereira et al., 2023), constitutions (Bai et al., 2022), or other rules set forth by the generative ghost’s creator. In addition to producing novel content, generative ghosts may also possess agentic (Park et al., 2023; Shavit et al., 2023) capabilities (i.e., the ability to autonomously execute complex sequences of actions with side effects in either the digital or physical realms).
@@
We expand on our definition of generative ghosts by discussing the design space of potential instantiations of this concept. While all generative ghosts share the capability to retrieve pre-existing multimodal content and generative novel multimodal content representing a deceased individual (including through the execution of agentic interactions), we argue that several key dimensions may impact the capabilities, perceptions, and societal impacts of such systems, including provenance, deployment timeline, anthropomorphism paradigm, multiplicity, cutoff date, embodiment, and representee type. As we detail in the Discussion, interrelationships between these dimensions are common. However, we present them separately here for the sake of clarity.
3.1. Provenance: First-Party vs. Third-Party Ghosts
@@
The provenance of a generative ghost refers to who created it (or authorized its creation). A first-party generative ghost is one created by the individual being represented (i.e., perhaps via end-of-life planning or legacy crafting practices). In contrast, a third-party generative ghost is not created by the representee. Third-party generative ghosts might be created by parties with a direct personal connection to the deceased (e.g., family, friends), by parties with a fiscal connection to the deceased (e.g., employers, estates), or by unconnected parties. Additionally, third-party ghosts may be authorized or unauthorized by the representee. Authorized third-party generative ghosts, for example, might be created with consent (e.g., via the deceased’s will). Meanwhile, unauthorized ghosts created by unconnected third parties are most likely in the case of public figures such as historical figures or contemporary celebrities; indeed, we already see early attempts at creating third-party generative ghosts of historical figures for entertainment and educational purposes by companies such as character.ai (Metz, 2023) and nonprofits such as Khan Academy (Grant, 2023).
3.2. Deployment Timeline: Pre- vs. Post-Mortem
@@
The deployment timeline of a generative ghost indicates whether the system’s initial deployment was pre-mortem or post-mortem. While some generative ghosts will be deployed post-mortem with the explicit purpose of memorializing the dead, many generative ghosts will likely start off as agents that serve some pre-mortem functionality (e.g., responding to personal emails or accomplishing other tasks on behalf of the representee, in character as if the representee had executed the task themself) – what we have termed generative clones – only to transition to ghosts after the death of their representee. Pre-mortem deployments might be desirable as they would allow an individual the opportunity to tune the behavior and/or capabilities of a first-party generative ghost to their liking, perhaps as part of an end-of-life planning process. Likewise, we imagine that generative clones might be designed with mortality in mind, potentially with attendant modifications to their behavior and/or capabilities once they become ghosts. Note that the ethical implications of pre-mortem generative clones (also called “digital twins,” “AI clones,” or “mimetic models”) have been discussed in related works such as (McIlroy-Young et al., 2022; Lee et al., 2023); our contribution focuses on identifying potential risks, benefits, and a research agenda around post-mortem clones, i.e., generative ghosts.
3.3. Anthropomorphism Paradigm: Reincarnations vs. Representations
@@
The anthropomorphism paradigm of a generative ghost refers to a subtle, but potentially impactful, interface metaphor choice: whether a generative ghost presents itself as a reincarnation of the deceased individual or as a representation of that individual. Generative ghosts employing a reincarnation metaphor would engage with end-users as if they were the deceased individual. In contrast, generative ghosts employing a representation metaphor would engage with end-users as if they are a representation of the deceased, but are not the deceased themselves. Design choices that might impact perceptions of whether a generative ghost is a reincarnation or a representation include whether a generative ghost uses first-person pronouns, whether it uses the present or past tense when discussing its representee, whether it uses the name of its representee or a different nomenclature (e.g., the use of “Fredbot” in (Kurzweil, 2023)), whether it is allowed to make statements that assert it is alive, possesses a soul, etc.
3.4. Multiplicity: Single Ghost vs. Multiple Ghosts
@@
The multiplicity design dimension refers to whether an individual is represented by a single generative ghost or by multiple distinct ghosts. Multiple ghosts might arise intentionally (i.e., a single creator develops a set of generative ghosts with different behaviors, capabilities, and/or audiences in mind). Multiple ghosts might enable people to customize different ghosts for different audiences in order to avoid forms of context collapse (Davis and Jurgenson, 2014; Litt, 2012; Marwick and Boyd, 2011). Multiple ghosts might also arise unintentionally due to multiple third parties creating generative ghosts for a single individual, or perhaps due to post-mortem identity theft or other novel cybercrimes.
3.5. Cutoff Date: Static vs. Evolving
@@
3.6. Embodiment: Physical vs. Virtual Ghosts
The embodiment dimension refers to whether a generative ghost has a physical embodiment. Such embodiments might be physical in a literal sense (e.g., robotics) or might be embodiments in rich digital media (e.g., avatars in mixed reality environments). In contrast, purely virtual generative ghosts would lack embodiment (e.g., perhaps available only as a chatbot). In addition to technical and/or cost constraints, there may be additional reasons to opt for lower-fidelity (i.e., virtual-only) embodiments, including potential ethical or psychological concerns related to physical embodiments.
3.7. Representee: Human vs. Non-Human Legacies
@@
4. Benefits and Risks of Generative Ghosts
@@
4.1. Generative Ghosts: Potential Benefits
@@
4.1.1. Potential Benefits for Representees
Individuals may choose to create first-party generative ghosts as part of their end-of-life planning for several reasons. On a personal level, people may feel comforted by or even feel hopeful in the idea of a “digital afterlife.” Prior work on digital legacy has documented people’s motivations to capture their life stories (Thomas and Briggs, 2014), major life events (Jamison-Powell et al., 2016), and pass on values to future generations (Gulotta et al., 2017). Posthumous, personalized AI representations may likewise be seen as a way for people to ensure that they are remembered, whether by family, friends, or perhaps the world at large.
@@ 4.1.1. Potential Benefits for Representees
4.1.2. Potential Benefits for the Bereaved
Interacting with a generative ghost may provide emotional support to close friends and family of the deceased. Prior research has considered the impact of online memorials on the bereaved (e.g. (Walter et al., 2012)), often responding to concerns that such spaces might prolong grief (see Section 4.2.1). Yet scholars cite the benefits of online memorials that allow loved ones to maintain healthy “continuing bonds” (Klass et al., 1996) with the deceased (e.g., (Getty et al., 2011; Brubaker and Hayes, 2011; Hjorth, 2021); c.f., (DeGroot, 2018)), often in a space where those grieving can gather (Carroll and Landry, 2010; Brubaker et al., 2013).
As with griefbots more generally, discussions with a generative ghost might provide comfort or closure and support a continued feeling of closeness. Generative ghosts may support positive mental health, particularly if they enable the bereaved to maintain connections with the deceased while “making sense” of the loss (termed “accommodation” (Neimeyer and Thompson, 2014)). Family members may find it comforting not only to interact with the generative ghost in the immediate post-mortem period, but to know that they might be able to interact with their deceased loved one at upcoming future life events such as weddings, the birth of a child, or other milestones during which they may wish to share updates with the generative ghost and/or hear advice, congratulations, or other event-appropriate commentary from their deceased loved one.
One-sided conversations with deceased loved ones, such as speaking graveside or writing to the deceased on a memorial, are common and well-recognized in the literature (DeGroot, 2018; Brubaker and Hayes, 2011; Brubaker et al., 2013) and are generally considered a part of healthy grieving (Klass et al., 1996). As summarized by DeGroot (DeGroot, 2018), these conversations are cathartic and help the living confront their loss: “People essentially struggle to develop an identity where the deceased is now part of the past self and not the present self… and communicating with the deceased appears to ease the transition” (DeGroot, 2018). However, it is worth noting that the dead do not typically respond. Rather, people imagine how the deceased would reply (DeGroot, 2018). As such, it is unclear how responses from a generative ghost might impact bereavement.
@@ 4.1.2. Potential Benefits for the Bereaved
4.1.3. Potential Benefits for Society
@@
4.2. Generative Ghosts: Potential Risks
@@
While thoughtful implementations of generative ghosts have potential upsides for the representee, their loved ones, and society more broadly, there are also risks. By proactively anticipating these risks, we may be able to design technical and sociotechnical systems that avoid or mitigate them. Likewise, we might be able to identify applications for which the risks are too high, regardless of thoughtful design. Of course, as with any powerful new technology, it is impossible to fully anticipate apriori the many potential sociotechnial side-effects of its introduction. The emergence of “AI afterlives” may reshape society in complex ways beyond our current imagining (i.e., spurring the advent of new religious movements). Here, we discuss four broad categories of risk associated with generative ghosts: mental health risks, reputational risks, security risks, and socio-cultural risks.
4.2.1. Mental Health Risks
@@
Earlier we discussed ways in which generative ghosts might offer mental health benefits to mourners. However, ghosts may also pose mental health risks for the bereaved. Scholars in thanatology and clinical psychology have long distinguished between adaptive coping – strategies that integrate the loss into an individual’s life, fostering emotional adjustment, well-being, and personal growth over time – and maladaptive coping – behavior that may obstruct healthy grieving, prolonging distress and resulting in depression, anxiety, and complicated grief (also known as Prolonged Grief Disorder)(Stroebe et al., 2013; Shear, 2015). Meanwhile, studies of prior generations of AI technologies (and even of non-AI computing systems (Reeves and Nass, 1996)) have found that people tend to anthropomorphize such systems and even develop attachments (Gabriel et al., 2024). As such, ghosts may include risks related to delayed accommodation, excessive attachment, information overload, anthropomorphism, deification, and second deaths. We anticipate that without careful design generative ghosts might magnify these risks due to the increasing fidelity of generative AI technologies and the modeling of the agent on a specific individual.
@@
Interacting with a generative ghost may impact the bereaved’s ability to assimilate and accommodate (Payne et al., 2007; Neimeyer and Thompson, 2014) their loss and integrate the reality of their loved one’s death (Klass et al., 1996). Delayed accommodation may negatively impact the bereaved’s mental health and quality of life by supporting loss-oriented experiences (e.g., active grieving, reminiscing while looking at old photos, etc.) at the expense of restorative-oriented experiences (e.g., the development of new relationships and life patterns) (Stroebe and Schut, 1999). Stroebe and Schut argue that both loss- and restorative-oriented experiences are necessary, but that having agency around when and how one oscillates between the two is important – agency that could be impeded by the interactive design of generative ghosts (e.g., if a generative ghost were designed to use “push notifications” that might pull a mourner into an interaction rather than letting the mourner decide when and how they want to engage in remembrance and other grief-related activities).
@@
A related concern is the possibility of developing excessive attachment to generative ghosts, as indicated by spending a disproportionate amount of time interacting with representations of the deceased in lieu of interacting with the living. Already, there are indications that some individuals find AI-powered companions, such as those from Replika (replika.ai), to be highly compelling (Liang, 2023); generative ghosts’ basis in real, beloved individuals rather than in fictional personas or historical figures may amplify this type of attachment.
@@
An additional mental health risk is that of second deaths or second losses, the experience of grieving a loved one again. [^1] Grief scholarship sometimes uses “secondary loss” to refer to experiences of grief that follow the “primary loss” of a loved one (e.g., loss of relationships, home, shared activities). Here, however, we focus on circumstances that might result in people engaging with their primary loss and grief a second time. In digital contexts, second deaths occur when data becomes unavailable either through technical obsolescence, deletion, or lack of access (Bassett, 2022). In the case of generative ghosts, second death refers to the emotional harm that may result if a ghost ceases to exist. Second deaths may occur for myriad reasons, including economic (i.e., the company that maintains the generative ghost service goes out of business; survivors’ or the deceased’s estate can no longer afford maintenance fees), regulatory (i.e., a government outlaws generative ghosts, which may be particularly likely in theocracies where AI afterlives threaten established belief systems), technological (i.e., obsolete generative ghosts no longer run on future infrastructure), or security reasons (i.e., a hacker disables a generative ghost). While second loss is not unique to AI systems, the potential intensity of engagement that generative ghosts can support means that the loss experienced at the disabling of a generative ghost might be especially profound.
4.2.2. Reputational Risks
@@
Privacy risks are context-dependent (Nissenbaum, 2009). For instance, it may be acceptable for a generative ghost to share sensitive information with some audiences, but not with others (e.g., perhaps it is acceptable for the ghost to discuss the deceased’s religious beliefs with their spouse, but not with their close work friend). The creation of multiple ghosts, each with different knowledge or abilities targeted toward different audiences, might mitigate privacy risks related to context collapse (Davis and Jurgenson, 2014; Litt, 2012; Marwick and Boyd, 2011).
@@
Fidelity risks refer to challenges that might arise from both accurate and inaccurate information. Related to hallucination risks, ghosts providing inaccurate information can be a problem, particularly in historical, legal, and economic applications. Yet accurate information also comes with risks. As with human memories, personal and family stories change over the years, and most forms of media decay over time. However, digital media defaults towards persistence, which can impede the important roles that forgetting (Mayer-Schonberger, 2009) and evolving memories can play (Pasupathi, 2001). Generative ghosts might disrupt the natural tendency for loved ones to recall the deceased through “rose-colored glasses” (i.e., the tendency to focus on positive memories and attributes) by persistently reminding mourners of particular negative events or character traits.
4.2.3. Security Risks